OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). cos x = sin x

OpenStudy (anonymous):

there are several ways to solve this problem

OpenStudy (anonymous):

@mathsminds would factoring work for this problem?

OpenStudy (anonymous):

\[\cos x = \sin x\]

OpenStudy (anonymous):

divide both sides by cos x

OpenStudy (anonymous):

\[\frac{ cosx }{ cosx }= \frac{ sinx }{ cosx }\]

OpenStudy (anonymous):

\[tanx=1\]

OpenStudy (anonymous):

\[x = \arctan(1)\]

OpenStudy (anonymous):

\[x=\frac{ \pi }{ 4 }\]

OpenStudy (anonymous):

@mathsmind and x = 5pi/4

OpenStudy (anonymous):

i want to show u how to get all the solutions between [0.2pi]

OpenStudy (anonymous):

if u know how to do that then all the best

OpenStudy (anonymous):

do u want to see another method of how to solve this question or not?

OpenStudy (anonymous):

ur solving for tan as being +ve

OpenStudy (anonymous):

@mathsmind no, thank you! those were the two solutions and I figured it out through your first method :)

OpenStudy (anonymous):

and u know that tan will be +ve in the first quarter and the third quarter is +ve

OpenStudy (anonymous):

but by the way u can't use that first method always, this is just a special case which u can do so, however u should get to the habit of using trig identities